V What is the pH
p
H
of a buffer system that contains 0.40 M
M
NH4+
N
H
4
+
and 0.20 M
M
NH3
N
H
3
? The pKa
p
K
a
of NH4+
N
H
4
+
is 9.25.
1
Expert's answer
2020-01-20T05:10:57-0500

Answer: According to the Henderson equation.

pH = pKa + log10([A-]/[HA])

Here [A-] is NH3

and [HA] is NH4+


Given [A-] = 0.2 M

[HA] = 0.4 M

Pka = 9.25


pH = pKa + log10([A-]/[HA])

= 9.25 + log10(0.2/0.4)

= 9.25 + log10(0.5)

= 9.25 - 0.30

= 8.95